2020 25 Staff Hourly Rates (changes on 1 Jan each year) Staff 2018 2019 Bar Staff A $ 21 $ 23 $ Bar Staff B $ 21 $ 23 $ Bar SDate 500 600 800 500 500 600 800 500 500 600 800 500 500 600 800 500 Product Profits per night Staff Pay Per night Day PayingTotal Product Profit Total Staff Cost Gross Nightly ProfitWhat are the formulas I need to find total product profit, total staff cost and gross nightly profit?
Answer #1

B10 fx =E2*0.8+F2*0.6+G2*0.5+H2*0.75+12*0.6+J2*0.5+K2*0.5+L2*0.5 B E L o R Nuts D Playing Patron Revenue Beer 100 5279870 141

11 =SUM(E2:E6) *0.8+SUM(F2:F6) *0.6+SUM(G2:56) *0.5+SUM(H2:H6) *0.75+SUM(12:16) *0.6+SUM(12:L6) *0.5 А B E L Q Date Day 05-07

M10 X =SUM(M2:02) B E L O Nuts D Playing Patron Revenue Beer 100 5279870 141 12053952 273 8355967 67 1165839 92 1020274 F Win

711 : =SUM(M2:06) А B E L 0 F Date D Playing Patron Revenue Beer 100 5279870 141 12053952 273 8355967 67 1165839 92 1020274 D

F15 3 X fx =D2+B10-M10 B E F L Q R K Chips Nuts 130 D Playing Patron Revenue Beer 100 5279870 141 12053952 273 8355967 67 116

16 fic А B E F L Q R Date Nuts =SUM(D2:06) +B11-M11 C D Playing Patron Revenue Beer 100 5279870 141 12053952 273 8355967 67 1

Notes

  • I used Randome numbers not the exact same numbers in the picture but the formula will be the same.
  • For the Gross Profit You can substract others costs too , You just have to choose the cell from the respected table.